at what temperature will aluminium have a resistivity that is three times resistivity copper has at room temperature?
1
Expert's answer
2020-04-06T09:14:19-0400

The temperature dependence of resistivity is given by equation

"\rho(t)=\rho(20\degree {\rm C})(1+\alpha (t-20\degree \rm C))."

Rresistivity of metals at room temperature

"{\rho_{Cu}(20\degree {\rm C})=1.68\times 10^{-8}\rm \Omega\cdot m}\\
\rho_{Al}(20\degree {\rm C})=2.65\times 10^{-8}\rm \Omega\cdot m\\"

Hence, we get equation

"3\times1.68\times 10^{-8} =2.65\times 10^{-8}(1+0.00435 (t-20\degree \rm C)),""{0.902 =0.00435 (t-20\degree \rm C)},\\
t=20+207=227\degree\rm C"